# “I think this is fair”: Uncovering the Complexities of Stakeholder Decision-Making in AI Fairness Assessment

### v1

@@ Abstract.
Assessing fairness in artificial intelligence (AI) typically involves AI experts who select protected features, fairness metrics, and set fairness thresholds. However, little is known about how stakeholders, particularly those affected by AI outcomes but lacking AI expertise, assess fairness. To address this gap, we conducted a qualitative study with 30 stakeholders without AI expertise, representing potential decision subjects in a credit rating scenario, to examine how they assess fairness when placed in the role of deciding on features with priority, metrics, and thresholds. We reveal that stakeholders’ fairness decisions are more complex than typical AI expert practices: they considered features far beyond legally protected features, tailored metrics for specific contexts, set diverse yet stricter fairness thresholds, and even preferred designing customized fairness. Our results extend the understanding of how stakeholders can meaningfully contribute to AI fairness governance and mitigation, underscoring the importance of incorporating stakeholders’ nuanced fairness judgments.
@@ 1. Introduction
Fairness in artificial intelligence (AI) has gained significant attention due to potential harms caused by relying on predicted outcomes, such as loan applications (Kurshan et al., 2022) and university admission (Marcinkowski et al., 2020). In developing AI models, their fairness has often been measured quantitatively through fairness metrics that evaluate whether a model treats individuals or groups in a fair manner (Verma and Rubin, 2018). Typically, AI experts, such as data scientists, AI developers, AI fairness researchers and practitioners, identify features that they would like to assess and protect, and then choose fairness metrics alongside thresholds that determine when outcomes are considered fair (Hort et al., 2024; Feldman et al., 2015; Saplicki and Bante, 2022). While this common expert-driven practice simplifies processes and achieves nominal technical AI fairness, it rarely aligns with human-centered fairness, which goes beyond AI expert perspectives and requires AI systems to reflect the values and priorities that diverse users and groups see as essential to fairness (Lee et al., 2017; Luo et al., 2025). This misalignment arises because people interpret fairness differently (Landers and Behrend, 2023). Recent studies further show that stakeholders without AI expertise (or non-experts), that is those without formal training or professional experience in AI model development, often hold fairness preferences that diverge from those of AI experts (Luo et al., 2025; Nakao et al., 2023). Moreover, even AI experts themselves report persistent challenges in determining which features to protect and which fairness metrics to apply (Madaio et al., 2022).
@@
We argue that these stakeholders, on equal footing with AI experts, need to be able to engage directly with questions of whom to protect (features), how to protect (metrics), and to what extent (thresholds). Such involvement not only increases transparency and fosters public trust but also contributes to actionable fairness strategies, provides crucial guidance to AI experts, and helps avoid the additional biases that may arise when stakeholders’ preferences are mediated.
@@
Our work pursues this aim, yet there remain four key research gaps that need to be addressed. First, current HCI and fairness practice led by AI experts may result in overly narrow definitions of fairness and introduce biases stemming from experts’ choices of features and metrics (Cheng et al., 2021). Indeed, to which features stakeholders pay attention has not been sufficiently investigated, even though previous research indicates that stakeholders have a broader fairness concerns, including features such as income in loan scenarios (Nakao et al., 2022). Second, current research does not capture how these features are prioritized by stakeholders to understand their fairness trade-offs. Third, current studies have attempted to solicit metric preferences from stakeholders (Alkhathlan et al., 2024; Wong, 2020; Cheng et al., 2021; Saxena et al., 2019; Haider et al., 2024; Ilvento, 2020; Yokota and Nakao, 2022) but primarily focus on soliciting stakeholders’ most preferred fairness metric and overlook how stakeholders perceive fairness nuances at the level of specific features (Cheng et al., 2021; Haider et al., 2024; Srivastava et al., 2019; Saxena et al., 2019). This limits opportunities for stakeholders to articulate their specific fairness needs (Nakao et al., 2023), and fails to uncover the complexity of how metric preferences interact with different features (Haider et al., 2024; Cheng et al., 2021). Last, only limited research has engaged stakeholders in deciding acceptable fairness thresholds (Luo et al., 2025), but those established by legal standards or determined by AI experts may again not necessarily align with stakeholder expectations. Therefore, general studies that begin with and focus solely on metric preferences (Saxena et al., 2019; Nakao et al., 2022; Srivastava et al., 2019; Haider et al., 2024; Alkhathlan et al., 2024; Luo et al., 2025; Cheng et al., 2021; Yu et al., 2020; Nakao et al., 2023) may offer AI experts some insight into how stakeholders define fairness. However, such an approach may not reveal which aspects of fairness stakeholders believe should be protected and prioritized, or whether the AI system is perceived as fair under the applied metrics, thereby possibly failing to align fairness with stakeholder values.
@@
To address these gaps, we reframe AI fairness assessment as a participatory process that empowers stakeholders to perform tasks traditionally reserved for AI experts. This process centers on three core lenses: (1) selecting and ranking features to determine whom to protect, (2) mapping features to fairness metrics to define how to protect them, and (3) setting thresholds to specify what level of unfairness is acceptable. We investigated this process through a qualitative study with 30 participants who lacked AI expertise but represented potential decision subjects. Each participant engaged in fairness assessment tasks through an interactive prototype system we designed to support this process. We chose a commonly used credit rating scenario to ground our investigation. By observing stakeholders perform fairness assessment, we show the complexity of their fairness decision-making while generating insights that can directly inform AI mitigation in real-world systems. We address the following research questions:
@@
  • RQ1 How do stakeholders identify features that are crucial for assessing AI fairness, and how do they prioritize them?
@@
  • Empirical evidence that stakeholders without AI expertise can conduct full AI fairness assessment, and that their practices reveal complexities substantially different from current expert-led approaches.
  • New insights into the importance of custom fairness metrics for stakeholders.
  • A detailed characterization of stakeholders’ feature, metric, and fairness thresholds choices observed in a credit rating scenario, yet offering patterns likely to generalize to other high-stakes AI domains.
  • Actionable design implications for interactive tools and governance processes that enable stakeholders to conduct full AI fairness assessment alongside experts, and leverage their insights to inform fairness mitigation in AI systems.
This paper is structured as follows. Section 2 reviews existing research and practices measuring AI fairness from an algorithmic perspective and stakeholder-centered approaches to AI fairness, identifying gaps our study addresses. Section 3 details our qualitative study in a credit rating scenario and the prototype system that operationalized expert-led fairness assessment for non-experts. Section 4 presents the results of each research question, and Section 5 reflects on stakeholder decision complexity, with design implications for future fairness tools.
@@ 2.1. Measuring AI Fairness
AI fairness has been conceptualized along three main perspectives: procedural fairness, i.e., fair processes leading to decisions (Robert et al., 2020; Lee et al., 2019), interactional fairness, i.e., focusing on fair interactions with individuals (Bies and Shapiro, 1987), and outcome fairness, i.e., evaluating whether the results of an AI system are fair (Rawls, 2001). Of these, outcome fairness is the most commonly assessed, typically using quantitative metrics. AI experts have proposed 109 different fairness metrics (Hort et al., 2024), with more than 20 widely used ones (Verma and Rubin, 2018), addressing different fairness needs to ensure equitable outcomes for individuals or groups (Pessach and Shmueli, 2022; Mehrabi et al., 2021; Hort et al., 2024). Of major importance in AI fairness are protected features, which are safeguarded by anti-discrimination laws (Hort et al., 2024; Mehrabi et al., 2021; Verma and Rubin, 2018), for example, protections against discrimination based on age, gender, and race under the United Kingdom’s Equality Act 2010 (UK Government, 2010), EU Law (European Union, 2012), or United States’ anti-discrimination laws in employment (Barocas and Selbst, 2016). Some sensitive features, such as “foreign worker” (Mehrabi et al., 2021; Kamiran and Calders, 2009; Friedler et al., 2019), while not strictly covered by law, are also sometimes deemed important in certain contexts as they may be related to protected characteristics such as race.
@@ 2.2. Stakeholders’ Perspectives of AI Fairness
In summary, while significant progress has been achieved, there is a need to investigate how to align technical AI fairness with human-centered notions of fairness; in short, we need to better understand how stakeholders perceive and assess AI fairness. Our work builds on this foundation by examining how stakeholders approach fairness decision-making, uncovering its complexities across feature selection and ranking, feature-metric pairing, and fairness threshold setting. By taking a human-centered perspective, we aim to support algorithmic fairness by incorporating the complex fairness needs of stakeholders.
@@ 3.1. Scenario Setup: Dataset and AI Model
Given that our participants were lay people without AI expertise, we carefully selected a scenario, a dataset and an AI model that would allow them to meaningfully engage with fairness assessment without imposing excessive cognitive load. To ground the study in a concrete and high-stakes context, we chose a credit rating scenario, a decision-making domain that is both familiar to the general public and consequential in real life.
Within this scenario, we employed the German Credit Dataset (Hofmann, 1994), one of the most frequently used datasets in fairness research, including HCI studies (Yang et al., 2024; Le Quy et al., 2022; Pagano et al., 2023; de Castro Vieira et al., 2025). This dataset presents a relatively simple binary classification task which classifies individuals as having “Good Credit” or “Bad Credit”. It contains 1,000 instances and 20 features, including multiple legally protected and unprotected features, enabling us to observe participants’ fairness assessment behavior while avoiding information overload. It includes two legally protected features (‘Age’ and ‘Gender’ (European Union, 2012)) and 18 unprotected features. ‘Age’ was converted into a binary feature, with individuals under 25 categorized as the protected group through thresholding at 25 (i.e., $\text{age}<25$ and $\text{age}\geq 25$ (Le Quy et al., 2022; Kamiran and Calders, 2009; Friedler et al., 2019)). ‘Gender’ was extracted from the ‘personal status and gender’ feature, designating ‘female’ as the protected group (Le Quy et al., 2022). Sometimes, ‘Foreign Worker’ is treated as a sensitive feature in the credit context (Kamiran and Calders, 2009; Friedler et al., 2019) because it can indirectly reveal legally protected features like race, but we did not specifically call it out in our study, leaving our participants to decide on its significance. The unprotected features were grouped into at most four distinct categories to avoid data sparsity, while numerical features were divided into up to four bins as needed.
We chose logistic regression as it is widely used in credit rating, where complex models add little accuracy (Hort et al., 2024). Its interpretability also helps participants provide fairness feedback (Barredo Arrieta et al., 2020; Cheng et al., 2021; Nakao et al., 2022). We trained the model using 5-fold cross-validation on all 1000 instances and selected one fold of 200 instances as our test dataset for user exploration. On the selected test dataset, we achieved a reasonable level of AI predictive performance with an accuracy of 0.76. Meanwhile, this test dataset can allow participants to explore diverse fairness assessment results across different features and metrics, ranging from 0% group differences (indicating no unfairness between groups) to 100% group differences (indicating maximum unfairness between groups).
@@ 3.2. Participant Recruitment
Our sample comprised 13 male and 17 female participants, with educational backgrounds ranging from secondary school (n=1) to bachelor’s degree (n=20), master’s degree (n=6), and doctoral degree (n=3). Ages ranged from 21 to 76: n=19 aged 20–30 years, n=8 aged 30–40 years, n=2 aged 40–50 years, and n=1 over 50. Our sample also encompassed varied professions and nationalities (Appendix A.1). We note that our participants were relatively young; however, this aligns with evidence from multiple countries showing high loan application activity among younger age groups (Average Age of First-Time Buyers in the UK: A Regional Breakdown, 2025; Average Student Loan Debt by Age 2025: Facts & Statistics, 2025; Personal Loan Statistics in Australia, 2025). Moreover, our sample balanced loan experience, with 14 participants reporting prior loan applications and 16 without such experience.
Each participant received the local equivalent of a \$19 Amazon voucher as compensation for their time and effort. We carefully considered the ethical implications of our study, particularly focusing on obtaining informed consent, privacy, and confidentiality. This study was reviewed and approved by University of Glasgow College of Science & Engineering Ethics Committee as a low-risk study (Application Number: 300230183).

### v2

# ”I think this is fair”: Uncovering the Complexities of Stakeholder Decision-Making in AI Fairness Assessment

@@ Abstract.
Assessing fairness in artificial intelligence (AI) typically involves AI experts who select protected features, fairness metrics, and set fairness thresholds to assess outcome fairness. However, little is known about how stakeholders, particularly those affected by AI outcomes but lacking AI expertise, assess fairness. To address this gap, we conducted a qualitative study with 26 stakeholders without AI expertise, representing potential decision subjects in a credit rating scenario, to examine how they assess fairness when placed in the role of deciding on features with priority, metrics, and thresholds. We reveal that stakeholders’ fairness decisions are more complex than typical AI expert practices: they considered features far beyond legally protected features, tailored metrics for specific contexts, set diverse yet stricter fairness thresholds, and even preferred designing customized fairness. Our results extend the understanding of how stakeholders can meaningfully contribute to AI fairness governance and mitigation, underscoring the importance of incorporating stakeholders’ nuanced fairness judgments.
@@
journalyear: 2026copyright: ccconference: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems; April 13–17, 2026; Barcelona, Spainbooktitle: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems (CHI ’26), April 13–17, 2026, Barcelona, Spaindoi: 10.1145/3772318.3790770isbn: 979-8-4007-2278-3/2026/04ccs: Human-centered computing Empirical studies in HCIccs: Computing methodologies Artificial intelligence
@@ 1. Introduction
Fairness in artificial intelligence (AI) has gained significant attention due to potential harms caused by relying on predicted outcomes, such as loan applications (Kurshan et al., 2022) and university admission (Marcinkowski et al., 2020). AI fairness has been conceptualized along three main perspectives: interactional fairness, i.e., focusing on fair interactions with individuals (Bies and Shapiro, 1987); procedural fairness, i.e., fair processes leading to an AI system’s outcomes (Robert et al., 2020; Lee et al., 2019); and outcome fairness, i.e., evaluating whether the outcomes of an AI system are fair (Rawls, 2001). Of these, outcome fairness is the most commonly assessed (Robert et al., 2020; Grgić-Hlača et al., 2018), and it is also the perspective we seek to explore in this work.
In practice, AI outcome fairness has often been measured quantitatively through fairness metrics that evaluate whether a model treats individuals or groups in a fair manner (Verma and Rubin, 2018). Typically, AI fairness researchers and practitioners with AI-related expertise whose task it is to design fair AI systems or mitigate existing AI systems so they are fair (Madaio et al., 2022)–who we refer to in this work as AI experts–identify features that they would like to assess and protect, and then choose fairness metrics alongside thresholds that determine when outcomes are considered fair (Hort et al., 2024; Feldman et al., 2015; Saplicki and Bante, 2022). While this common expert-driven practice simplifies processes and achieves nominal technical AI fairness, it rarely aligns with human-centered fairness, which goes beyond AI expert perspectives and requires AI systems to reflect the values and priorities that diverse individuals and groups see as essential to fairness (Lee et al., 2017; Luo et al., 2025). This misalignment arises because people interpret fairness differently (Landers and Behrend, 2023). Recent studies further show that stakeholders without AI expertise, that is those without formal training or professional experience in AI model development, often differ from AI experts in defining AI fairness (Luo et al., 2025; Nakao et al., 2023). Moreover, even AI experts themselves report persistent challenges in determining which features to protect and which fairness metrics to apply (Madaio et al., 2022).
@@
We argue that these stakeholders on equal footing with AI experts, need to be able to engage directly with questions of whom to protect (features), how to protect (metrics), and to what extent (thresholds). Such involvement not only increases transparency and fosters public trust but also contributes to actionable fairness strategies, provides crucial guidance to AI experts, and helps avoid the additional biases that may arise when stakeholders’ preferences are mediated.
@@
Our work pursues this aim by supporting and uncovering how those who are affected by AI outcomes and also lack AI expertise (whom we refer to as lay stakeholders, or simply stakeholders in this work) assess AI outcome fairness. However, there remain four key research gaps that need to be addressed. First, current HCI and fairness practice led by AI experts may result in overly narrow definitions of fairness and introduce biases stemming from experts’ choices of features and metrics (Cheng et al., 2021). Indeed, to which features stakeholders pay attention has not been sufficiently investigated, even though previous research indicates that stakeholders have a broader fairness concerns, including features such as income in loan scenarios (Nakao et al., 2022). Second, current research does not capture how these features are prioritized by stakeholders to understand their fairness trade-offs. Third, current studies have attempted to solicit metric preferences from stakeholders (Alkhathlan et al., 2024; Wong, 2020; Cheng et al., 2021; Saxena et al., 2019; Haider et al., 2024; Ilvento, 2020; Yokota and Nakao, 2022) but primarily focus on soliciting stakeholders’ most preferred fairness metric and overlook how stakeholders perceive fairness nuances at the level of specific features (Cheng et al., 2021; Haider et al., 2024; Srivastava et al., 2019; Saxena et al., 2019). This limits opportunities for stakeholders to articulate their specific fairness needs (Nakao et al., 2023), and fails to uncover the complexity of how metric preferences interact with different features (Haider et al., 2024; Cheng et al., 2021). Last, only limited research has engaged stakeholders in deciding acceptable fairness thresholds (Luo et al., 2025), but those established by legal standards or determined by AI experts may again not necessarily align with stakeholder expectations. Therefore, general studies that begin with and focus solely on metric preferences (Saxena et al., 2019; Nakao et al., 2022; Srivastava et al., 2019; Haider et al., 2024; Alkhathlan et al., 2024; Luo et al., 2025; Cheng et al., 2021; Yu et al., 2020; Nakao et al., 2023) may offer AI experts some insight into how stakeholders define fairness. However, such an approach may not reveal which aspects of fairness stakeholders believe should be protected and prioritized, or whether the AI system is perceived as fair under the applied metrics, thereby possibly failing to align fairness with stakeholder values.
@@
To address these gaps, we reframe AI fairness assessment as a participatory process that empowers stakeholders to perform tasks traditionally reserved for AI experts. This process centers on three core lenses: (1) selecting and ranking features to determine whom to protect, (2) mapping features to fairness metrics to define how to protect them, and (3) setting thresholds to specify what level of unfairness is acceptable. We investigated this process through a qualitative study with 26 participants who lacked AI expertise but represented potential decision subjects. Each participant engaged in fairness assessment tasks through an interactive prototype system we designed to support this process. We chose a commonly used credit rating scenario to ground our investigation. By observing stakeholders perform fairness assessment, we show the complexity of their fairness decision-making while generating insights that can directly inform AI mitigation in real-world systems. We address the following research questions:
@@
  • RQ1 How do stakeholders identify features that are crucial for assessing AI outcome fairness, and how do they prioritize them?
@@
  • We provide empirical evidence showing how stakeholders without AI expertise assess AI outcome fairness across features, metrics, and fairness thresholds in a credit-rating context. Our findings shed light on the complexity of their reasoning and the generic patterns that differ from common expert practices.
  • We offer new insights into the importance of unprotected features and custom fairness metrics for stakeholders.
  • We derive actionable design implications for interactive tools and governance processes that make AI fairness assessment more accessible to stakeholders, enable stakeholders to conduct AI fairness assessment alongside experts, and leverage their insights to inform fairness mitigation in AI systems.
This paper is structured as follows. Section 2 reviews existing research and practices measuring AI fairness from an algorithmic perspective and stakeholder-centered approaches to AI fairness, identifying gaps our study addresses. Section 3 details our qualitative study in a credit rating scenario and the prototype system that operationalized outcome fairness assessment for lay stakeholders. Section 4 presents the results of each research question, and Section 5 reflects on stakeholder decision complexity, with design implications for future fairness tools.
@@ 2.1. Measuring AI Fairness
Among the three perspectives on AI fairness, interactional fairness is relatively hard to measure, since it considers whether people understand the information an AI system provides and feel respected in the interaction (Robert et al., 2020). Procedural fairness has gained attention, though standardized metrics remain limited; prior works often examine which features should be used, whether inappropriate or sensitive features (e.g., race, gender) are included, or ways to restrict their use (Belitz et al., 2021; Grgić-Hlača et al., 2018). Outcome fairness, in comparison, is the perspective most commonly translated into quantitative assessments and AI experts have proposed 109 different fairness metrics (Hort et al., 2024), with more than 20 widely used ones (Verma and Rubin, 2018), addressing different fairness needs to ensure equitable outcomes for individuals or groups (Pessach and Shmueli, 2022; Mehrabi et al., 2021; Hort et al., 2024).
Of major importance in AI outcome fairness are protected features, which are safeguarded by anti-discrimination laws (Hort et al., 2024; Mehrabi et al., 2021; Verma and Rubin, 2018; Rabonato and Berton, 2025), for example, protections against discrimination based on age, gender, and race under the United Kingdom’s Equality Act 2010 (UK Government, 2010), EU Law (European Union, 2012), or United States’ anti-discrimination laws in employment (Barocas and Selbst, 2016). Some sensitive features, such as “foreign worker” (Mehrabi et al., 2021; Kamiran and Calders, 2009; Friedler et al., 2019), while not strictly covered by law, are also sometimes deemed important in certain contexts as they may be related to protected characteristics such as race.
@@ 2.2. Stakeholders’ Perspectives of AI Fairness
In summary, while significant progress has been achieved, there is a need to investigate how to align technical AI fairness with human-centered notions of fairness; in short, we need to better understand how stakeholders perceive and assess AI outcome fairness. Building on this foundation, our work empowers stakeholders to perform fairness assessment tasks in ways similar to AI experts, and we examine how stakeholders approach fairness decision-making, uncovering its complexities across feature selection and ranking, feature-metric pairing, and fairness threshold setting. By taking a human-centered perspective, we aim to support algorithmic fairness by incorporating the complex fairness needs of stakeholders.
@@ 3.1. Scenario Setup: Dataset and AI Model
Given that our participants were lay stakeholders, we carefully selected a scenario, a dataset, and an AI model that would allow them to meaningfully engage with fairness assessment without imposing excessive cognitive load. To ground the study in a concrete, high-stakes context, we chose a credit rating scenario, a decision-making domain familiar to the public and consequential in real life.
Within this scenario, we employed the German Credit Dataset (Hofmann,Hans, 1994), one of the most frequently used datasets in fairness research (Le Quy et al., 2022; de Castro Vieira et al., 2025; Pagano et al., 2023), including HCI studies (Yang et al., 2024; Luo et al., 2025) where lay stakeholders need the ability to reflect on fairness in an understandable setting. This dataset presents a relatively simple binary classification task which classifies individuals as having “Good Credit” or “Bad Credit”. It contains 1,000 instances and 20 features, including multiple legally protected and unprotected features, enabling us to observe participants’ fairness assessment behavior while avoiding information overload. It includes two legally protected features (‘Age’ and ‘Gender’ (European Union, 2012)) and 18 unprotected features. ‘Age’ was converted into a binary feature, with individuals under 25 categorized as the protected group through thresholding at 25 (i.e., $\text{age}<25$ and $\text{age}\geq 25$ (Le Quy et al., 2022; Kamiran and Calders, 2009; Friedler et al., 2019)). ‘Gender’ was extracted from the ‘personal status and gender’ feature, designating ‘female’ as the protected group (Le Quy et al., 2022). Sometimes, ‘Foreign Worker’ is treated as a sensitive feature in the credit context (Kamiran and Calders, 2009; Friedler et al., 2019) because it can indirectly reveal legally protected features like race, but we did not specifically call it out in our study, leaving our participants to decide on its significance. The unprotected features were grouped into at most four distinct categories to avoid data sparsity, while numerical features were divided into up to four bins as needed. Appendix A.1 provides detailed descriptions of each of the 20 features that participants could refer to, based on the original dataset documentation (Hofmann,Hans, 1994; Le Quy et al., 2022).
We chose logistic regression as it is widely used in credit rating, where complex models add little accuracy (Hort et al., 2024). Its interpretability also helps participants provide fairness feedback (Barredo Arrieta et al., 2020; Cheng et al., 2021; Nakao et al., 2022). We trained the model using 5-fold cross-validation on all 1000 instances and selected one fold of 200 instances as our test dataset for user exploration. On the selected test dataset, we achieved a reasonable level of AI predictive performance with an accuracy of 0.76.
Meanwhile, the test dataset can allow participants to explore diverse fairness assessment results across different features and metrics, ranging from 0% group differences (indicating no unfairness between groups) to 100% group differences (indicating maximum unfairness between groups).
@@ 3.2. Participant Recruitment
We conducted the study locally through in-person sessions. While this approach allowed us to closely observe participants’ decisions and reasoning, we recognize its regional limitation and therefore adopted purposive sampling, to prioritize diversity in nationality (resulting in 10 different nationalities) and profession (resulting in 11 different types), with educational backgrounds ranging from secondary school (n=1) to bachelor’s degree (n=20), master’s degree (n=6), and doctoral degree (n=3), considering that different lived experiences and cultural backgrounds might influence people’ views on fairness. Then, we tried to balance gender (13 male and 17 female participants) and age. Ages ranged from 21 to 76 (19 aged 20–30 years, 8 aged 30–40 years, 2 aged 40–50 years, and 1 over 50). We note that our participants were relatively young; however, this aligns with evidence from multiple countries showing high loan application activity among younger age groups (Average Age of First-Time Buyers in the UK: A Regional Breakdown, 2025, 2025, 2025). Moreover, our sample balanced loan experience, with 14 participants reporting prior loan applications and 16 without such experience.
It is important to note that AI fairness can be perceived and conceptualized from different perspectives. Our study focuses on uncovering how participants make outcome fairness assessments, a perspective that aligns with current popular auditing practice by AI experts: whom to protect (features),how to protect them (metrics), and to what extent (thresholds). However, we cannot assume that all participants hold or prefer to assess fairness through this perspective. Indeed, among the 30 participants we recruited, we did observe differences in preferred fairness perspectives. We accounted for this in our data analysis (Section 3.5). Full participant information is available in Table 4 in Appendix A.2.
Every one of the 30 recruited participants received a £15 Amazon voucher as compensation for their time and effort. We carefully considered the ethical implications of our study, particularly focusing on obtaining informed consent, privacy, and confidentiality. This study was reviewed and approved by the University of Glasgow College of Science & Engineering Ethics Committee (Application Number: 300230183) as a low-risk study.